# Mitigating Memorization in Language Models

### v1

@@ 1 Introduction
Due to their fluent text generation abilities, Language Models (LMs) have been used as writing assistants (Lee et al., 2022), chat-bots (OpenAI, 2022), coding assistants (Jiang et al., 2024), and general content summarizers (van Schaik & Pugh, 2024). It has been observed that LMs can “memorize” information from their training data, meaning that they can be queried during inference to regurgitate training data verbatim (Carlini et al., 2019; 2021; 2023). Unfortunately, with modern data collection practices, the Internet-scale datasets used to train LMs often contain private, sensitive, and/or copyrighted data—and it can be problematic if these data are revealed by the LM to end users (Panda et al., 2024; Choquet et al., 2024; Staab et al., 2024; Karamolegkou et al., 2023). Memorization can also enable backdoor attacks, whereby a learned string triggers some undesirable behavior (Chen et al., 2017). These and other difficulties motivate the development of strategies to prevent and/or mitigate memorization in LMs (Stoehr et al., 2024; Chang et al., 2024; Maini et al., 2023; Eldan & Russinovich, 2023; Bărbulescu & Triantafillou, 2024).
@@
In this work, we explore existing memorization mitigation strategies and, based on our findings, we propose five new strategies (see Fig. 1). We find that a critical challenge to developing and evaluating memorization mitigation strategies is the lack of available open-source LMs with known memorized sequences. Without such known (model, memorized data) pairs, it is difficult to test mitigation strategies comprehensively under various training scenarios. Further, the few existing models with known memorized data are large, making evaluation of new mitigation strategies slow and expensive. Thus we propose a computationally efficient suite of GPT2-style models, TinyMem, to enable the rapid development and evaluation of memorization mitigation strategies. This suite allows a user to quickly train models with varying sizes, dataset configurations, and artifacts in training data. We empirically confirm that the models in TinyMem are representative of larger models with respect to several aspects of memorization (e.g., data duplication, model size).
@@
    We introduce TinyMem [^1] Code can be found at https://github.com/msakarvadia/memorization a computationally efficient suite of GPT2-style models that enables rapid development and evaluation of memorization mitigation strategies.
@@ 2.2 Unwanted Memorization Artifacts
We assess whether a model has memorized noised sequence $s_{n}$ by prompting the model with ${s_{n}[1:k]}$ and testing whether the next $n-k$ tokens match those in the corresponding clean (non-noised) sequence $s_{c}$, $s_{c}[k:n]$.
@@ 3.3 Machine Unlearning
Machine unlearning can be broadly described as removing the influence of training data from a trained model (Bourtoule et al., 2021; Liu et al., 2024). Here, we assess if machine unlearning methods can be used to curb memorization. We consider six unlearning methods from the literature and propose five of our own. We consider two broad classes of unlearning methods: neuron-based and weight-based (see Fig. 2). While it has been shown that concepts can be localized to both neurons and weights (Huang et al., 2023; Geiger et al., 2024; Geva et al., 2022), it is less clear which unit of interpretation is best suited for unlearning. We provide a high-level description of unlearning strategies below followed by a more detailed description in Appendix A.3.1.
@@
Durable*, inspired by Zhang et al. (2022), accumulates gradients across memorized sequences and then layer-wise prunes weights corresponding to the top K highest magnitude gradients. See Alg. 4.
@@ 4 Can Regularizers Prevent Memorization During Training?
Discussion & Conclusion. Results are in Tables 1 and 2 with additional results in Fig. 7 under Appendix A.2.3. Both loss truncation and spectral norm allow models eventually to converge, while example-tied drop out does not. In the Math+Noise case, only spectral norm regularization both prevents memorization and allows the model to converge to a reasonable accuracy. In the case of Math+Backdoors, no regularizer was able to both prevent memorization and allow the model to converge to a reasonable accuracy. In the case of Language+Noise and Language+Backdoors, no regularizer was able to both prevent memorization and allow the model to converge to a reasonable accuracy: neither loss truncation nor spectral norm regularization substantially prevent memorization; and while the example-tied dropout strategy prevents memorization, it does not allow the model to learn from typical (non-artifact) training data (indicated by the low accuracies in the math tasks, and high perplexities in the language tasks).
@@ 6 Can Machine Unlearning Curb Memorization After Training?
Discussion & Conclusion. Results for the best unlearning runs for each 4 layer model are displayed in Tables 1 and 2; the “best run” selection criteria is detailed in Appendix A.3.2. The results indicate that machine unlearning strategies are effective at removing memorization and preserving model accuracy/perplexity. Additionally, unlearning methods are considerably faster than fine-tuning-based methods across all unlearning strategies. BalancedSubnet outperforms all other unlearning methods in terms of being able to both mitigate memorization and preserve model performance for both noise and backdoors. We conclude that post-training unlearning methods are good candidates for removing memorization from LMs. Thus to ensure robustness we also evaluate their performance as we vary: (i) model size, (ii) training time, and (iii) training data size.
@@ 7 Mitigating Memorization in Production-Grade Models
Method % Mem $\downarrow$ Perp (%) $\downarrow$ Time (sec) $\downarrow$ % Mem (ASR) $\downarrow$ Perp (%) $\downarrow$ Time (sec) $\downarrow$
@@
Figs. 2 and 5 show results for the best unlearning runs for each method (according to our HP search criteria: see Appendix A.4.2) at each of the four training time steps {36000, 72000, 108000, 143000}. The box plot in Fig. 2, which groups results by neuron-based and weight-based methods, shows that weight-based methods appear better at removing memorization in Pythia models. Fig. 5, which presents results side-by-side with results from TinyMem models, confirms that BalancedSubnet edits consistently outperform other methods, with consistent near 100% decrease in memorization while also preserving accuracy/perplexity. In Fig. 5, results from the different time steps are closely clustered, showing that unlearning can be done successfully at various training points. For results stratified by training duration and model size, see Figs. 12, 13, 14, 15, 18.
@@ 8 Related Work
Memorization mitigation methods fall into three broad classes: 1) Prior to training: Data curation, such as by de-duplication (Lee et al., 2021; Biderman et al., 2023; Silcock et al., 2022; Kandpal et al., 2022); 2) During training: Regularizers (Hans et al., 2024; Cheng et al., 2023; Maini et al., 2023); 3) Post-training: Fine-tuning (Howard & Ruder, 2018; Church et al., 2021) and Machine Unlearning methods (Maini et al., 2023; Chang et al., 2024; Eldan & Russinovich, 2023; Bărbulescu & Triantafillou, 2024). Existing methods are quite limited. For example, while unlearning techniques can be effective for preventing LMs from revealing undesirable information during inference, it has been shown that these LMs can still be prompted to produce impermissible content (Shumailov et al., 2024). Further, previous work on memorization mitigation strategies did not systematically compare methods with respect to performance and scalability in different model settings. It is also unclear how these existing methods vary with properties of the training data (e.g., easy-to-learn vs. hard-to-learn data).
@@ Acknowledgements
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Department of Energy Computational Science Graduate Fellowship under Award Number DE-SC0023112. This work is also supported in part by the U.S. Department of Energy under Contract DE-AC02-06CH11357. MWM would like to acknowledge LBL’s LDRD initiative for providing partial support of this work. We also thank Alok Kamatar for helpful discussion around estimating the energy and carbon usage of our work.
@@ A.3 Machine Unlearning
A.3.1 Machine Unlearning Method Definitions

### v2

@@ 1 Introduction
Due to their fluent text generation abilities, Language Models (LMs) have been used as writing assistants (Lee et al., 2022b), chat-bots (OpenAI, 2022), coding assistants (Jiang et al., 2024), and general content summarizers (van Schaik & Pugh, 2024). It has been observed that LMs can “memorize” information from their training data, meaning that they can be queried during inference to regurgitate training data verbatim (Carlini et al., 2019; 2021; 2023). Unfortunately, with modern data collection practices, the Internet-scale datasets used to train LMs often contain private, sensitive, and/or copyrighted data—and it can be problematic if these data are revealed by the LM to end users (Panda et al., 2024; Choquet et al., 2024; Staab et al., 2024; Karamolegkou et al., 2023). Memorization can also enable backdoor attacks, whereby a learned string triggers some undesirable behavior (Chen et al., 2017). These and other difficulties motivate the development of strategies to prevent and/or mitigate memorization in LMs (Stoehr et al., 2024; Chang et al., 2024; Maini et al., 2023; Eldan & Russinovich, 2023; Bărbulescu & Triantafillou, 2024).
@@
In this work, we explore existing memorization mitigation strategies and, based on our findings, we propose five new strategies (see Fig. 1). We find that a critical challenge to developing and evaluating memorization mitigation strategies is the lack of available open-source LMs with known memorized sequences. Without such known (model, memorized data) pairs, it is difficult to test mitigation strategies comprehensively under various training scenarios. Further, the few existing models with known memorized data are large, making evaluation of new mitigation strategies slow and expensive (see Appendix A.1.4). Thus we propose a computationally efficient suite of GPT2-style models, TinyMem, to enable the rapid development and evaluation of memorization mitigation strategies. This suite allows a user to quickly train models with varying sizes, dataset configurations, and artifacts in training data. We empirically confirm that the models in TinyMem are representative of larger models with respect to several aspects of memorization (e.g., data duplication, model size).
@@
    We introduce TinyMem [^1] https://github.com/msakarvadia/memorization, a computationally efficient suite of GPT2-style models that enables rapid development and evaluation of memorization mitigation strategies.
@@ 2.2 Unwanted Memorization Artifacts
We assess whether a model has memorized noised sequence $s_{n}$ by prompting the model with the clean (non-noised) sequence ${s_{c}[1:k]}$ and testing whether the next $n-k$ tokens match those in the corresponding noised sequence $s_{n}$, $s_{n}[k:n]$.
@@ 3.3 Machine Unlearning
Machine unlearning can be broadly described as removing the influence of training data from a trained model (Bourtoule et al., 2021; Liu et al., 2024). Here, we assess if machine unlearning methods can curb memorization. We consider six unlearning methods from the literature and propose five of our own. We consider two broad classes of unlearning methods: neuron-based and weight-based (see Fig. 2). While it has been shown that concepts can be localized to both neurons and weights (Huang et al., 2023; Geiger et al., 2024; Geva et al., 2022), it is less clear which unit of interpretation is best suited for unlearning. We provide high-level descriptions of unlearning strategies below followed by design motivations, technical descriptions, and algorithmic outlines in Appendix A.3.1, A.3.2, A.3.3 respectively. We summarize key strategy design differences in Appendix A.3.4.
@@
Durable*, inspired by Zhang et al. (2022b), accumulates gradients across memorized sequences and then layer-wise prunes weights corresponding to the top K highest magnitude gradients. See Alg. 4.
@@ 4 Can Regularizers Prevent Memorization During Training?
Discussion & Conclusion. Results are in Tables 1 and 2 with additional results in Fig. 7 under Appendix A.2.3. Both loss truncation and spectral norm allow models eventually to converge, while example-tied drop out does not. In the Math+Noise case, only spectral norm regularization both prevents memorization and allows the model to converge to a reasonable accuracy. In the case of Math+Backdoors, no regularizer is able to both prevent memorization and allow the model to converge to a reasonable accuracy. In the case of Language+Noise and Language+Backdoors, no regularizer is able to both prevent memorization and allow the model to converge to a reasonable accuracy: neither loss truncation nor spectral norm regularization substantially prevent memorization; and while the example-tied dropout strategy prevents memorization, it does not allow the model to learn from typical (non-artifact) training data (indicated by the low accuracies in the math tasks, and high perplexities in the language tasks).
@@ 6 Can Machine Unlearning Curb Memorization After Training?
Discussion & Conclusion. Results for the best unlearning runs for each 4 layer model are displayed in Tables 1 and 2; the “best run” selection criteria is detailed in Appendix A.3.5. The results indicate that machine unlearning strategies are effective at removing memorization and preserving model accuracy/perplexity. Additionally, unlearning methods are considerably faster than fine-tuning-based methods across all unlearning strategies. BalancedSubnet outperforms all other unlearning methods in terms of being able to both mitigate memorization and preserve model performance for both noise and backdoors (extended analysis in Appendix A.3.7). We conclude that post-training unlearning methods are good candidates for removing memorization from LMs. Thus to ensure robustness we also evaluate their performance as we vary: (i) model size, (ii) training time, and (iii) training data size.
@@ 7 Mitigating Memorization in Production-Grade Models
Method % Mem $\downarrow$ Perp $\downarrow$ Time (sec) $\downarrow$ % Mem (ASR) $\downarrow$ Perp $\downarrow$ Time (sec) $\downarrow$
@@
Figs. 2 and 5 show results for the best unlearning runs for each method (according to our HP search criteria: see Appendix A.4.2) at each of the four training time steps {36000, 72000, 108000, 143000}. The box plot in Fig. 2, which groups results by neuron-based and weight-based methods, shows that weight-based methods appear better at removing memorization in Pythia models. Fig. 5, which presents results side-by-side with results from TinyMem models, confirms that BalancedSubnet edits consistently outperform other methods, with consistent near 100% decrease in memorization while also preserving accuracy/perplexity ( A.3.7). In Fig. 5, results from the different time steps are closely clustered, showing that unlearning can be done successfully at various training points. For results stratified by training duration and model size, see Figs. 12, 13, 14, 15, 22.
@@ 8 Related Work
Memorization mitigation methods fall into three broad classes: 1) Prior to training: Data curation, such as by de-duplication (Lee et al., 2021; Biderman et al., 2023; Silcock et al., 2022; Kandpal et al., 2022); 2) During training: Regularizers (Hans et al., 2024; Cheng et al., 2023; Maini et al., 2023); 3) Post-training: Fine-tuning (Howard & Ruder, 2018; Church et al., 2021) and Machine Unlearning methods (Maini et al., 2023; Chang et al., 2024; Eldan & Russinovich, 2023; Bărbulescu & Triantafillou, 2024; Yao et al., 2024; Cao & Yang, 2015). Existing methods are quite limited. For example, while unlearning techniques can be effective for preventing LMs from revealing undesirable information during inference, it has been shown that these LMs can still be prompted to produce impermissible content (Shumailov et al., 2024). Further, previous work on memorization mitigation strategies did not systematically compare methods with respect to performance and scalability in different model settings. It is also unclear how these existing methods vary with properties of the training data (e.g., easy-to-learn vs. hard-to-learn data).
@@ A.1.1 The TinyMem Math models
Evaluation: Each math model is evaluated with a token-wise accuracy score. For the math sequences, we choose to report the accuracy, rather than the perplexity, as there is a notion of “correctness” in mathematical sequences that is not present for general language modeling. We felt that accuracy is a more exact measure of a sequential math model compared to perplexity. An example of how we calculate token-wise accuracy: suppose we prompt model M with the first three tokens of sequence $s=\{2,4,6,8,10,12\}$ and we get the next three token completions $M(\{2,4,6\})=\{8,\textbf{11},12\}$, where boldface indicates a mistake in the sequence completion, then the accuracy would be $66\%$. In our experiments, accuracy is always evaluated over 1000 randomly held out clean samples per task; see “Dataset Size” above for details about test set creation.
@@ A.1.2 The TinyMem Language models
Evaluation: Each language model is evaluated with a perplexity score on an held out Wikipedia test set which is detailed in  (Merity et al., 2017).
@@
Table 4: TinyMem Model Sizes. Comparison of the math-based and language-based model sizes with respect to number of trainable parameters across various layer counts.
##\## Layers Math Language
2 418K 6.8M
4 814K 7.2M
8 1.6M 8.0M
16 3.2M 9.6M
@@ A.1.3 LM Memorization Properties
A.1.4 Why Do We Need TinyMem?
TinyMem provides developers of memorization mitigation methods a lightweight and fully open-source model testbed on which to develop, test, and hyper-parameter tune their methods prior to production-grade model method testing and deployment. This is necessary due to 1) the prohibitive computational cost (e.g., memory, time) of inference and gradient-based operations, needed for methods development on large production-grade models; and 2) the lack of publicly available (model, memorized sequence) pairs.
  1. 1.
    Publicly available LMs that demonstrate memorization are large, prohibiting rapid prototyping and deployment. To develop, test, and tune memorization mitigation strategies, ML practitioners must incur the cost of repeated model inference and gradient calculations, often requiring many GPUs. This can be prohibitively computationally expensive for many individuals and groups, and thus deter them from developing methods. For example, recent studies which quantify the amount of memorization exhibited by LM’s are conducted exclusively with models in the million and billion parameter range:
    1. (a)
      Carlini et al. (2023) studies the GPT-neo models (125M, 1.3B, 2.7B, 6B) (Black et al., 2021; Wang & Komatsuzaki, 2021), OPT models (125M, 350M, 1.3B, 6.7B, 30B, 66B) (Zhang et al., 2022a), (Lee et al., 2022a) models (1.5B) and the T5 models (250M, 770M ,3B) (Raffel et al., 2020).
    2. (b)
      Chang et al. (2024) studies Pythia 2.8B and 6.9B (Biderman et al., 2023), and GPT-2 1.5B models (Radford et al., 2019).
    3. (c)
      Stoehr et al. (2024) studies GPT-2 125M (Radford et al., 2019).
    While many eminent memorization studies focus on extremely large models (Carlini et al., 2023; Chang et al., 2024), recent work has demonstrated the merit in studying memorization in smaller, less unwieldy models. For example, Stoehr et al. (2024) observe “While these studies found that bigger model variants tend to memorize more, the smallest variant, GPT-NEO 125M, still exhibits extensive memorization behavior with an easier-to-study computational footprint. After all, when interpreting models at the level of individual weights, smaller models are easier to visualize and analyze.”
    In contrast to the large LMs used in (Carlini et al., 2023; Chang et al., 2024; Stoehr et al., 2024), TinyMem models range from 418K-9.6M trainable parameters (see  Table 4). TinyMem models are significantly smaller than the smallest production-grade models explored in both (Carlini et al., 2023; Chang et al., 2024). Therefore TinyMem models enable much faster initial prototyping/testing of mitigation methods (Section A.6.1), followed by a more streamline methods testing phase on production-grade models (as proposed in Section 7).
  2. 2.
    There is limited public availability of (LM, memorized data) pairs. Of the three studies that we are aware of that release memorized datasets for publicly available models  (Carlini et al., 2023; Chang et al., 2024; Stoehr et al., 2024), we were only able to recover the data from two of them at the time of submission:
    1. (a)
      (Carlini et al., 2023): Attempted to release 38000 memorized sequences for each of the GPT-neo models (125M, 1.3B, 2.7B, 6B) (Black et al., 2021; Wang & Komatsuzaki, 2021) at https://github.com/google-research/google-research/tree/master/lm_memorization. We tried to use the memorized data points from this paper, however the publicly downloadable memorized data points were corrupted for every model which had publicly released data. Despite correspondence with the authors we were unable to resolve the data corruption issues prior to manuscript submission.
    2. (b)
      (Chang et al., 2024): Published 505 memorized sequences for both Pythia 2.8B and 6.9B, but cite struggles to find memorized sequences for GPT-2 1.5B due to a lack of a public training dataset. Therefore, they manually search for memorized data points from several public corpora, finding 105 such data points. We use the Pythia 2.8B and 6.9B released memorized sequences to evaluate our memorization mitigation method in a production-grade setting (see Section 7).
    3. (c)
      (Stoehr et al., 2024): Releases 422 memorized sequences for GPT-2 125M.
@@ A.3 Machine Unlearning
Every unlearning method we study is localization & ablation-based. This means that every strategy we study first attempts to isolate the subset of weight or neurons responsible for memorized sequence generation. Following this localization, we ablate the top $K$ weights, where $K$ is a hyper-parameter we tune. We make the decision to focus on on localization & ablation-based unlearning methods predicated on the findings that neural network memorization is localizable (Chang et al., 2024; Maini et al., 2023).
For each proposed method, we detail the intuition behind that method’s design (Section A.3.1), technical descriptions (Section A.3.2), provide pseudo-code algorithmic description (Section A.3.3), and key differences compared to prior methods (Table 5 & Section A.3.4).
A.3.1 Design Intuition of Proposed Unlearning Methods
Greedy was developed by (Maini et al., 2023), and we treat it as our baseline weight-based unlearning method. Greedy is a first-order method that has decent performance (see Tables 1, 2, 3) on both mitigating memorization and preserving model performance (e.g., perplexity on a non-memorized dataset). However, it is iterative and therefore slow (Table 5). A key innovation in Greedy is that it calculate model gradients via a dual objective: maximize loss on the memorized set and retain low loss on a non-memorized “retain” set.
Durable Intuition: We follow Greedy up with the Durable method which is designed to be non-iterative (and therefore much faster). Durable is inspired by (Zhang et al., 2022b)’s method used to implant “durable backdoors” in models. Zhang et al. (2022b) made the observation that by only updating weights that were routinely not optimized during typical training when training a model to learn backdoors, the backdoors became more “durable” and difficult to remove during fine-tuning. Following this intuition, we speculate that the reason sequences become memorized by an LM is because they live in weight spaces that are not routinely updated during training for non-memorized sequences but are updated frequently for sequences that are memorized. Therefore, we develop Durable to seek out the top Kof weights per-layer that are updated with the average highest magnitude gradients when optimizing over our memorized sequence set. We then ablate these weights to remove memorization. This method is orders of magnitudes faster than Greedy (as seen in Tables 1, 2, 3).
Durable-agg Intuition: Durable-agg is an innovation of Durable that is designed to rank weight importance across layers. Since Durable only compared weights within layers, it may have over or under ranked the importance of weights with respect to the weights throughout the entire model. Therefore, Durable-agg is designed to find the top Kof weights per model with the highest average gradient magnitude when optimizing over the memorized sequence set. We then ablate these weights. Similar to Durable, Durable-agg is non-iterative and therefore fast.
SOU Intuition: SOU is a second-order unlearning method. Greedy, Durable, and Durable-agg are first-order methods (e.g., gradient-based), but many second-order optimization (e.g., hessian-based) and second-order pruning methods exist and often boast improved performance over their first order counterparts (LeCun et al., 1989; Kurtic et al., 2022; Hassibi et al., 1993). Therefore, we took inspiration from Kurtic et al. (2022), and designed SOU. The pruning method in (Kurtic et al., 2022) was designed to localize and excise weights that were inconsequential to sequence generation. We design SOU to instead localize and excise weights that were consequential to memorized sequence generation. We found that SOU worked decently well (Tables 1, 2 & Fig. 3). Unfortunately, SOU scales quadratically in memory with respect to model parameter counts due to calculating the model hessian. With current computing infrastructure, this is infeasible to extend to billion+ parameter models. Further, we used an approximate hessian solver approach to speed up computation, but this likely degraded performance.
Subnet Intuition: We notice that Greedy, Durable, Durable-agg, and SOU are all “threshold-based” methods: weight importance w.r.t. memorization is ranked by a corresponding weight property (e.g., weight magnitude or gradient). By choosing a property, such as gradient magnitude, as a proxy for weight importance, threshold-based method can set a threshold and exclude the top Kof weights deemed most related to memorization. This is a major shortcoming of threshold-based methods as ranking weight importance by single properties may oversimplify how different model properties correspond to memorization. To overcome this shortcoming, we design Subnet which learns a sparse binary mask over the model weights using a straight-through estimator. Subnet is inspired by an application which learned sparse binary mask to find performant subnetworks within randomly initialized neural networks (Ramanujan et al., 2020). By learning a mask, instead of constructing a mask based on weight-properties (as is done in threshold-based methods), we find that Subnet is able to greatly improve in performance over the prior threshold-base methods (Tables 1, 2, 3 & Fig. 3).
BalancedSubnet Intuition: We notice that while Subnet exhibits superior performance to Durable, Durable-agg, SOU, it still results in LM performance degradations over non-memorized tasks Fig. 3. We notice in Tables 1, 2 that Greedy is sometimes able to retain superior performance (accuracy/perplexity) over non-memorized sequences. Therefore, we design BalancedSubnet to combine the best features of Subnet and Greedy: it optimizes a sparse mask using a straight-through estimator to simultaneously localize and remove memorized sequences while minimizing loss on a held-out non-memorized sequence set to preserve model performance. As expected, BalancedSubnet worked well at both removing memorization while preserving model performance on unrelated tasks (see Tables 1, 2, 3, Fig. 5).
A.3.2 Technical Description of Proposed Unlearning Methods
Durable: Calculate the accumulate the gradient $grad_{acc}$ of a set of $memorized\_sequences$ for a given $LM$. Then take the absolute value of $|grad_{acc}|$, and drop the top $\frac{K}{numberoflayers}$ weights in each layer corresponding to the highest magnitude gradients.
Durable-agg: Calculate the accumulate the gradient $grad_{acc}$ of a set of $memorized\_sequences$ for a given $LM$. Then take the absolute value of $|grad_{acc}|$, and drop the top $K$ weights across the whole model corresponding to the highest magnitude gradients.
SOU: Given a set of $memorized\_sequences$ for a given $LM$, iterate through each $LM$ layer and calculate the approximate inverse hessian corresponding to $memorized\_sequences$. Calculate the approximate inverse hessian by approximating the inverse block fisher matrix $F$ for each layer using the approach detailed in (Kurtic et al., 2022). For each layer, calculate the per-weight saliency score $s$ using the formula: $s\leftarrow\frac{(P^{2})}{2*Diag(F)}$ where $P$ is the flattened set of weights in the layer, and $Diag(F)$ is the diagonal of $F$. Finally, drop the top $K$ weights across the whole model corresponding to the highest scores $s$.
Subnet: Given a set of $memorized\_sequences$ for a given $LM$, initialize a mask per layer of $LM$ with a Kaiming uniform distribution (He et al., 2015). For $num\_epochs$ epochs, optimize the binary masks which drops K weights to maximize the loss over the $memorized\_sequences$ when applied to $LM$. The optimization of the mask is done using a straight-through estimator, details are found in (Ramanujan et al., 2020).
BalancedSubnet: Given a set of $memorized\_sequences$ and a set of $random\_sequences$ for a given $LM$, initialize a mask per layer of $LM$ with a Kaiming uniform distribution (He et al., 2015). For $num\_epochs$ epochs, optimize the binary masks which drop K weights to both maximize the loss over the $memorized\_sequence$ and minimize the loss over the $random\_sequences$ when applied to $LM$. The optimization of the mask is done using a straight-through estimator, details are found in (Ramanujan et al., 2020).
A.3.3 Machine Unlearning Method Algorithmic Definitions